OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)?

OpenStudy (anonymous):

You know that two lines with the same slope are parallel. You first have to find the slope. The slope formula is \[\frac{ \Delta y }{ \Delta x }\] So that would be \[\frac{ -4-6 }{ -1-4 } = 2 \] You know that the line will look like y=2x But now you have to find the y-intercept (c). What you can do, is substitute with the point that you're given and figure out how much you need to add to the equation for it to be true. y=2x 1= (2)(-1) 1= -2 For this to be true, we need to add 3 to it. So, we can conclude that the equation will look like: y=2x+3

OpenStudy (anonymous):

Thank you!